# AI & Human Co-Improvement for Safer Co-Superintelligence

### v1

@@ Abstract
AI that improves itself has been the primary goal of the field since its inception [1]. Historically, practical instantiations have focused on parameterizing a model in terms of weights, and then finding the best choice of those weights, without human intervention – from linear models through to neural networks. The 2010s onwards marked the era of scaling to ever-larger models [2], yielding large performance gains, but still focused on self-improving via weights only, while the architecture, data, objective function, update rule, and implementation (code) were mostly fixed. The current era has expanded that self-improvement search to improving all aspects through learning: models that create their own training data [3], challenge themselves to be better [4, 5], and learn to evaluate and reward themselves based on their own performance on these tasks [6, 7, 8]; see Table 2. Some of these axes have already provided sizable gains in performance, and approaches like synthetic data creation and LLM-as-a-Judge are now standard building blocks for frontier models. The quest for AI that improves its own architectures and rewrites its own code is still in its infancy [9, 10], but early signs show its promise, with a current push on autonomous AI research agents [11, 12, 13, 14].
It seems clear by now that we are marching towards ever more intelligent AI systems that in the long run will surpass humans in all task metrics, and by a large margin. Fully realized self-improvement is clearly an end-game marker. However, endowing AIs with this autonomous ability without appropriate guidance built into the system is fraught with danger for humankind – from misuse through to misalignment [15]. Nevertheless, there is still time left before AI eclipses humans in all endeavors, and in AI research in particular. We thus propose that it is better to focus on humans and AI working together to solve these problems.
@@
There are multiple plausible pathways to superintelligence, and current AI systems are already helping us explore these research directions to some degree, e.g. through code assistance and writing assistance. And, in general, improving the overall capabilities of frontier models does equip them with some of the skills suitable for research collaboration – as a byproduct. However, we typically see that skills improve further when they are targeted, for example lots of effort has been spent on improving AI coding skills, resulting in improved AI coding ability. However, while coding isn’t solved either [16], there is much more to “solving” AI than coding alone. Our central point is that with more development resources spent on endowing AI with AI research collaboration skills, these abilities will get better too.
@@
Our position is related to treatise on human-centered AI [46, 47, 48, 49], but in our case more specifically for the goal of achieving (co-)superintelligence via collaborative research. Similarly, the work of [50] advocates for cooperative AI, and finding common ground with machines, where research would be a special case. Nevertheless, we agree that work on all forms of human-AI collaboration is also an important goal for future human society with the advent of superhuman AIs. Other works emphasize that misalignment is a challenging goal, and advocate for the importance of AI’s objectives to be human-oriented [51]. Importantly, our position is that this challenging problem can and should be tackled by collaboration. For example, collaboration with AI may more readily help find flaws in their own designs; however, we believe this research should be in progress now – not added on later when systems are fully deployed and it is too late.
@@
Various works have described autonomous self-improvement and possible ways it can be achieved, for example [9, 52, 53, 54, 55]. Correspondingly, a large number of works have also developed practical instantiations with various technical contributions, see Table 2 for examples.

### v2

@@ Abstract
AI that improves itself has been the primary goal of the field since its inception [1]. Historically, practical instantiations have focused on parameterizing a model in terms of weights, and then finding the best choice of those weights, without human intervention – from linear models through to neural networks. The 2010s onwards marked the era of scaling to ever-larger models [2], yielding large performance gains, but still focused on self-improving via weights only, while the architecture, data, objective function, update rule, and implementation (code) were mostly fixed. The current era has expanded that self-improvement search to improving all aspects through learning: models that create their own training data [3], challenge themselves to be better [4, 5], and learn to evaluate and reward themselves based on their own performance on these tasks [6, 7, 8]; see Table 2. Some of these axes have already provided sizable gains in performance, and approaches like synthetic data creation and LLM-as-a-Judge are now standard building blocks for frontier models. The quest for AI that improves its own architectures and rewrites its own code is still in its infancy [9, 10], but early signs show its promise, with a current push on autonomous AI research agents [11, 12, 13, 14, 15].
It seems clear by now that we are marching towards ever more intelligent AI systems that in the long run will surpass humans in all task metrics, and by a large margin. Fully realized self-improvement is clearly an end-game marker. However, endowing AIs with this autonomous ability without appropriate guidance built into the system is fraught with danger for humankind – from misuse through to misalignment [16, 17]. Nevertheless, there is still time left before AI eclipses humans in all endeavors, and in AI research in particular. We thus propose that it is better to focus on humans and AI working together to solve these problems.
@@
There are multiple plausible pathways to superintelligence, and current AI systems are already helping us explore these research directions to some degree, e.g. through code assistance and writing assistance [18, 19, 20, 21]. And, in general, improving the overall capabilities of frontier models does equip them with some of the skills suitable for research collaboration – as a byproduct [22, 23]. However, we typically see that skills improve further when they are targeted, for example lots of effort has been spent on improving AI coding skills, resulting in improved AI coding ability. However, while coding isn’t solved either [24], there is much more to “solving” AI than coding alone. Our central point is that with more development resources spent on endowing AI with AI research collaboration skills, these abilities will get better too.
@@
Our position is related to treatise on human-centered AI [54, 55, 56, 57], but in our case more specifically for the goal of achieving (co-)superintelligence via collaborative research. Similarly, the work of [58] advocates for cooperative AI, and finding common ground with machines, where research would be a special case. The recent AI co-scientist [59] is a related practical system, but for the biomedical domain, not for AI research. Nevertheless, we agree that work on all forms of human-AI collaboration is also an important goal for future human society with the advent of superhuman AIs. Other works emphasize that misalignment is a challenging goal, and advocate for the importance of AI’s objectives to be human-oriented [60]. Scalable oversight is a related solution of allowing humans to intervene in the alignment process by providing supervision [61]. Importantly, our position is that this challenging problem can and should be tackled by research collaboration, which is bidirectional. For example, collaboration with AI may more readily help find flaws in their own designs, which also shares common goals with [62]. Overall, we believe this research should be in progress now – not added on later when systems are fully deployed and it is too late.
@@
Various works have described autonomous self-improvement and possible ways it can be achieved, for example [9, 63, 64, 65, 66]. Frontier labs are clearly aiming to develop self-improving AI, whilst also knowing that it has serious long-term risks [17, 67, 68]. Correspondingly, a large number of works have also developed practical instantiations with various technical contributions, see Table 2 for examples.